1.A bill for $21900 drawn on July 10 for 6 months was discounted for $21720 at 5% p.a.On what date was the bill discounted.
2.If the B.G. on a bill 146days hence at 6% p.a. is $360,find T.D.&B.D.
3.If the T.D. on a bill of $660 is $110.The rate of interest is 4%.In how much time will the bill be due?
4.Find B.D. and discounted value of a bill worth $600 drawn on May 15,2005 for 3 months and discounted on july 20, 2005 at 5%p.a.
1
Expert's answer
2020-06-02T20:18:02-0400

1.

21900-21720=180

"180=21900\times\frac{5}{100}\times n"


"n=\frac{180\times100}{21900\times5}=0.164383562"

"0.164383562\times365=60" days

The bill was drawn on July 10 for 6 months so, legal due date is 10

January. The bill was discounted 60 days before thus counting backwards from 10 January:

January - 10 days

December - 31 days

November - 19 days

The bill was discounted 11 November

2.

"360=TD\times\frac{146}{365}\times\frac{6}{100}"

TD=15000

BG=BD-TD

360=BD-15000, BD=15360

3.

BV=660

TD=110

PV=BV-TD=660-110=550

"110=550\times n\times\frac{4}{100}"

100=220n

n=5

4.

Bill is drawn on 15th May so legally due date is 15th 

August. As bill is discounted on 20th July.

Remaining period= 11days of July+15days of August=26

So, 

"BD=600\times\frac{26}{365}\times\frac{5}{100}=2.14"

The holder will receive =(600-2.14)=597.86